Find the local and absolute extreme values of the function on the given interval. Also

specify the intervals where function is increasing or decreasing

(i) 𝑓(π‘₯) = (π‘₯2+ π‘₯ + 1)2

1
Expert's answer
2022-01-10T19:09:30-0500

"f'(x)=\frac{\mathrm{d}}{\mathrm{d} x}\left[\left(x^{2}+x+1\right)^{2}\right]\\
f'(x)=2\left(x^{2}+x+1\right) \cdot \frac{\mathrm{d}}{\mathrm{d} x}\left[x^{2}+x+1\right]\\
f'(x)=2\left(x^{2}+x+1\right)\left(\frac{\mathrm{d}}{\mathrm{d} x}\left[x^{2}\right]+\frac{\mathrm{d}}{\mathrm{d} x}[x]+\frac{\mathrm{d}}{\mathrm{d} x}[1]\right)\\
f'(x)=2\left(x^{2}+x+1\right)(2 x+1+0)\\
f'(x)=2(2 x+1)\left(x^{2}+x+1\right)"


At "f'(x)=0,"


"2(2 x+1)\left(x^{2}+x+1\right)=0"

which implies,


"x=-\frac{1}{2}"

Put the value of "x" into the given function, we have:

"\text{Local Minimum } = (x,f(x)) = (-\frac{1}{2},\frac{9}{16})"


The function has no absolute extreme value.


The interval where the function is increasing or decreasing is thus:

"\text { Decreasing: }-\infty<x<-\frac{1}{2}\\ \text { Increasing: }-\frac{1}{2}<x<\infty"